OpenStudy (devonhoward15):

Complete the table by providing justification for each step of the proof. Conjecture: If 4(x^2-2)=28, then xis 3 or -3

OpenStudy (devonhoward15):

@Vocaloid

Vocaloid (vocaloid):

may I see the table?

OpenStudy (devonhoward15):

im trying to get it

OpenStudy (devonhoward15):

@Vocaloid

Vocaloid (vocaloid):

any ideas on the firs step?

OpenStudy (devonhoward15):

plug in x

Vocaloid (vocaloid):

no, i mean on the table

OpenStudy (devonhoward15):

no idea

Vocaloid (vocaloid):

note: the first step on the table is just the problem copied down, so we write "given"

Vocaloid (vocaloid):

any ideas on the second row?

OpenStudy (devonhoward15):

ohh so its just words describing what they are?

Vocaloid (vocaloid):

yes, it's words describing how to go from the previous step to the next step

OpenStudy (devonhoward15):

ohh ok

OpenStudy (devonhoward15):

is 2 simplified?

Vocaloid (vocaloid):

so for step 2, what rule do we use to go from 4(x^2 - 2) to 4x^2 - 8?

Vocaloid (vocaloid):

hint: it's called the ________ Property what word goes into the blank?

OpenStudy (devonhoward15):

distributive?

Vocaloid (vocaloid):

very good!

Vocaloid (vocaloid):

so for step 2 we can write distributive property

Vocaloid (vocaloid):

any ideas for step 3?

Vocaloid (vocaloid):

hint: it's called the ______ Postulate

OpenStudy (devonhoward15):

ohh i wont know that one

Vocaloid (vocaloid):

when we add the same thing to both sides, we call it the Addition Postulate

Vocaloid (vocaloid):

to go from 4x^2 - 8 = 28 to 4x^2 = 36, we add 8 to both sides

Vocaloid (vocaloid):

with me so far?

OpenStudy (devonhoward15):

yes

Vocaloid (vocaloid):

there are also multiplication, division, subtraction postulates that work the same way

OpenStudy (devonhoward15):

so addition postulate?

Vocaloid (vocaloid):

good! that's step 3

Vocaloid (vocaloid):

any ideas on 4?

OpenStudy (devonhoward15):

simplified

Vocaloid (vocaloid):

not quite

Vocaloid (vocaloid):

what do we do to go from 4x^2 = 36 to x^2 = 9?

Vocaloid (vocaloid):

is it addition, subtraction, division, or multiplication?

OpenStudy (devonhoward15):

division

Vocaloid (vocaloid):

good, so it's the division postulate

Vocaloid (vocaloid):

step 5 would just be "taking the square root" I don't think it has a special name

OpenStudy (devonhoward15):

how about simplified?

Vocaloid (vocaloid):

i'm not sure

Vocaloid (vocaloid):

"simplify by taking the square root?"

Vocaloid (vocaloid):

hm

OpenStudy (devonhoward15):

ill put that... teacher wont mind